(cos(3x)+isin(3x))/(cos(x)+isin(x))


1
Expert's answer
2021-10-21T11:47:12-0400

Given expression is: "\frac{(cos(3x)+isin(3x))}{(cos(x)+isin(x))}"

Using Euler form, this can simplify to: "\frac{e^{i(3x)}}{e^{i(x)}}\\"

"=e^{i(2x)}\\
=cos(2x)+isin(2x)"